# StreakNet-Arch: An Anti-scattering Network-based Architecture for Underwater Carrier LiDAR-Radar Imaging

### v2

@@ Abstract
In this paper, we introduce StreakNet-Arch, a novel signal processing architecture designed for Underwater Carrier LiDAR-Radar (UCLR) imaging systems, to address the limitations in scatter suppression and real-time imaging. StreakNet-Arch formulates the signal processing as a real-time, end-to-end binary classification task, enabling real-time image acquisition. To achieve this, we leverage Self-Attention networks and propose a novel Double Branch Cross Attention (DBC-Attention) mechanism that surpasses the performance of traditional methods. Furthermore, we present a method for embedding streak-tube camera images into attention networks, effectively acting as a learned bandpass filter. To facilitate further research, we contribute a publicly available streak-tube camera image dataset. The dataset contains 2,695,168 real-world underwater 3D point cloud data. These advancements significantly improve UCLR capabilities, enhancing its performance and applicability in underwater imaging tasks. The source code and dataset can be found at https://github.com/BestAnHongjun/StreakNet.
@@ I Introduction
Underwater laser imaging signal processing technology is crucial for obtaining underwater images, including 2D gray-scale maps and 3D point clouds images, which has wide applications in ocean exploration, biology [1], surveillance [2], archaeology, unmanned underwater vehicles control [3, 4], etc. In contrast to image processing algorithms for underwater image enhancement [5, 6, 7, 8, 9, 10, 11, 12, 13] or restoration [14, 15, 16], underwater laser imaging signal processing technology can process signals from a more fundamental source, such as streak-tube cameras and ICCD cameras. This approach enables the achievement of superior spatial resolution and extended detection ranges. However, its effectiveness is significantly hindered by a major challenge: scattering. This phenomenon drastically reduces image clarity and limits imaging range.
To address this, the Underwater Carrier LiDAR-Radar (UCLR) employs a suite of strategies to suppress scattering and achieve long-distance underwater imaging [17, 18, 19, 20]. Specifically, the UCLR’s laser source typically utilizes blue or green light to minimize propagation attenuation in water [21], thereby enhancing detection distance. Additionally, a range-gated detector is employed for the UCLR, which is sensitive only to reflected signals received within a specific time window after the pulse is emitted. More importantly, lasers are modulated into high-frequency pulses to exceed the cut-off frequency of water’s low-pass response [22, 23], effectively suppressing light scattering. Since the frequency is typically high ($\geq$100 MHz), receivers employing high temporal resolution optical detection devices are required, such as nanosecond-resolution ICCD cameras [24] or picosecond-resolution streak-tube cameras [25]. Underwater laser imaging relies on signal processing algorithms to extract target echoes from the received signal. These algorithms determine the presence and arrival time of the echoes, ultimately reconstructing the image. The processing typically involves two stages: scatter suppression and echo identification. Scatter suppression methods in the UCLR include bandpass filtering [25], adaptive filtering [26, 27, 28, 29, 30, 31], and machine learning-based filtering [32]. The objective is to process a signal containing scatter noise into a suppressed scatter signal. Echo identification methods in UCLR primarily rely on thresholding techniques, including manually set thresholds [25] and adaptive thresholds [33], often coupled with matched filtering approaches [25]. These methods aim to determine the presence of echo signals in the received signal.
@@
Figure 1: Overview of the StreakNet-Arch based UCLR system.
In this paper, we experimented with employing Self-Attention mechanism networks [35] in the signal processing phase of the UCLR to improve the scatter resistance performance. This network architecture, achieving state-of-the-art (SOTA) performance in diverse fields like computer vision [36, 37, 38, 39, 40] and natural language processing [35, 41, 42], emerges as a powerful candidate for a universal model. To prevent network overfitting to the signal itself and enhance the network’s generalization ability across different scenes, we also provide a template signal alongside the input signal. This guides the network to discern the presence of echo signals by learning the differences between the signal and the template signal. To this end, we have made adjustments to the Self-Attention mechanism, introducing a Double Branch Cross Attention (DBC-Attention) mechanism, which our experiments have proven to be more suited for underwater imaging tasks. Furthermore, addressing the limitations of existing algorithms in real-time imaging, we depart from the conventional two-stage processing paradigm. Instead, we formulate the task as an end-to-end binary classification. This approach directly determines whether an input signal contains echoes, thereby equipping the UCLR with the capability for real-time imaging. Given that our scene dataset is acquired using a streak-tube camera, and the signals extracted from captured streak images serve as inputs for the network, we have named this end-to-end architecture StreakNet-Arch.
@@
    We enhance the UCLR’s signal processing with Self-Attention networks. Further, we propose DBC-Attention, a groundbreaking variant specifically optimized for underwater imaging tasks. Experimental results conclusively demonstrate DBC-Attention’s superiority over the standard Self-Attention approach.
@@ III-A StreakNet-Arch
The proposed StreakNet-Arch based UCLR system (Fig. 1a-d) employs a sub-nanosecond Q-switch laser to generate subcarrier-modulated pulses at a frequency of 500 MHz with 532 nm. A portion of the generated pulse passes through the beam splitter into a delay device, which can be gated by a delay of $t_{G}$ seconds.  [^1] Range-gated imaging technology, which captures images by controlling the camera shutter delay for a certain period $t_{G}$. This enables the reception of signals within a specific range, mitigating the impact of backscattering on imaging. After the delay, a trigger signal is sent to the control circuit of the streak-tube camera. Simultaneously, another part of the pulse is reflected into the water tank.
@@ III-C Attention Analysis Method
(a) Neuronal perspective.
(b) Input perspective.
Figure 2: Different perspectives of MP model.
@@ III-G Datasets
(a) 25 m experimental water tank.
(b) 30 cm diameter experimental target.
Figure 3: Experimental setup.
(a) 20 m train
(b) 20 m valid
(c) 20 m test
(d) 15 m train
(e) 15 m valid
(f) 15 m test
(g) 13 m train
(h) 13 m valid
(i) 13 m test
(j) 10 m train
(k) 10 m valid
(l) 10 m test
Figure 4: Schematic diagram of dataset partitioning.
@@ IV-A Model training
TABLE II: The parameters of model training.
Model name Backbone architecture $\lambda_{w}$ Number of layers Number of heads
StreakNet-s Self-Attention 0.125 1 2
StreakNet-m Self-Attention 0.250 2 4
StreakNet-l Self-Attention 0.500 4 8
StreakNet-x Self-Attention 1.000 8 16
StreakNetv2-s DBC-Attention 0.125 1 2
StreakNetv2-m DBC-Attention 0.250 2 4
StreakNetv2-l DBC-Attention 0.500 4 8
StreakNetv2-x DBC-Attention 1.000 8 16
Under the StreakNet-Arch, we simultaneously trained StreakNet with the Self-Attention mechanism as the backbone network and StreakNetv2 with the DWC-Attention mechanism for comparative experiments.
During the training phase, the Stochastic Gradient Descent (SGD) algorithm is used to optimize for 120 epochs, with a base learning rate of $2\times 10^{-6}$ per batch. A cosine annealing learning rate strategy is employed (Fig. 5), and the Exponential Moving Average (EMA) method is used. The training was performed on a single NVIDIA RTX 3090 (24G). The parameters of all trained models are shown in Table II. After training, the loss decreased by approximately 70% (Fig. 6).
Figure 5: Learning rate variation curve during the training phase.
Figure 6: Loss variation curve during the training phase.
Figure 7: $F_{1}$ scores evaluated on the validation set for traditional algorithms and StreakNet.
@@ IV-B StreakNet-Arch exhibits superior anti-scattering capabilities compared to traditional imaging methods
We evaluated the traditional algorithms and the StreakNets on the validation set, and the results are shown in Fig. 7 and Table III. Clearly, both Self-Attention-based StreakNet and DWC-Attention-based StreakNetv2 achieve higher $F_{1}$ scores than traditional algorithms. This demonstrates that the StreakNet-Arch has stronger anti-scattering capabilities compared to traditional algorithms. The imaging results are shown in Fig. 8 and 9.
(a) Traditional 20 m
(b) StreakNet 20 m
(c) StreakNetv2 20 m
(d) Traditional 15 m
(e) StreakNet 15 m
(f) StreakNetv2 15 m
(g) Traditional 13 m
(h) StreakNet 13 m
(i) StreakNetv2 13 m
(j) Traditional 10 m
(k) StreakNet 10 m
(l) StreakNetv2 10 m
Figure 8: 2D imaging experimental results under various distances for each method.
(a) Traditional 20 m
(b) StreakNet 20 m
(c) StreakNetv2 20 m
(d) Traditional 15 m
(e) StreakNet 15 m
(f) StreakNetv2 15 m
(g) Traditional 13 m
(h) StreakNet 13 m
(i) StreakNetv2 13 m
(j) Traditional 10 m
(k) StreakNet 10 m
(l) StreakNetv2 10 m
Figure 9: 3D imaging experimental results under various distances for each method.
TABLE III: $F_{1}$ scores (%) for traditional imaging methods and StreakNet-Arch imaging methods on the validation set.
Model Traditional (baseline) StreakNet StreakNetv2
s 70.82 86.78 +15.96 86.92 +16.10
m 70.82 88.23 +17.41 87.03 +16.21
l 70.82 86.71 +15.89 86.35 +15.92
x 70.82 85.57 +14.75 86.33 +15.51
Figure 10: The sequence chart of traditional imaging algorithm.
$\text{AIT}_{\text{traditional}}=\frac{N+1}{2}t_{m}=\frac{N+1}{2}t_{m}.$ (22)
To measure the real-time imaging capability of the algorithm, we propose an evaluation metric called the Average Imaging Time (AIT), defined as the average time from the input of a streak-tube image to obtaining the corresponding imaging result. If we assume $t_{11}=t_{21}=...=t_{N1}=t_{m}$, the AIT for traditional imaging algorithms is calculated using Eq. 22.
Figure 11: The sequence chart of StreakNet-Arch method.
$\text{AIT}_{\text{streaknet}}=\frac{1}{N}\sum_{i=1}^{N}t_{i1}=t_{m}.$ (23)
It is evident that the AIT of the traditional imaging algorithm is a linear function with respect to $N$, while the AIT of StreakNet-Arch is a constant. Therefore, theoretically, in practical scenarios where $N$ is large, StreakNet-Arch will have a significant advantage in real-time imaging. To validate this theory, we conducted a comparative experiment. We sequentially input $N$ steak-tube images, where $N$ gradually increases from 1 to 64, and tested the AIT metric for both traditional algorithms and StreakNet-Arch on an NVIDIA RTX 3060 (12G) GPU. Experimental results are depicted in Fig. 12 and Table IV, with AIT values measured in milliseconds (ms). The experimental findings indicate that as the number of streak-tube images increases from 2 to 64, the AIT for traditional imaging methods escalates linearly from 58 ms to 1257 ms. In contrast, the AIT for the StreakNet method remains constant within the range of 54 ms to 84 ms.
Figure 12: Curve of AIT (ms) with the changing number $N$ of streak-tube images.
TABLE IV: AIT (ms) for traditional imaging algorithms and StreakNet-Arch algorithms.
@@ IV-D FD Embedding Layer is an equivalent bandpass filter
Since the carrier frequency of the detection signal is 500 MHz (see Fig. 1e), traditional imaging algorithms use a handcraft bandpass filter with a range of 450 MHz - 550 MHz during filtering. If we consider the bandpass filter from the perspective of “attention distribution”, we can think of the bandpass filter as a binary attention distribution with values of 1 for frequencies in the range of 450 MHz - 550 MHz and 0 for frequencies outside this range. The FD Embedding Layer’s attention distribution offers a similar concept, functioning as a learnable generalized bandpass filter.
@@
(a) Attention distribution with frequency after AAM analysis for StreakNet-m.
(b) Attention distribution with frequency after AAM analysis for StreakNetv2-m.
Figure 13: Results of attention distribution with frequency after AAM analysis.
Therefore, we further enumerated the range of bandpass filters in the range of 0 - 200 MHz, with each group spanning 5 MHz, and used traditional methods for imaging. We then calculated the $F_{1}$ score on the validation set. The experimental results are shown as the red curve in Fig. 14a. A peak appears at 42.5 MHz (i.e., the 40 MHz - 45 MHz bandpass range), indicating that frequency information near 40 MHz is indeed strongly correlated with anti-scattering imaging.
@@
(a) The $F_{1}$ scores on the validation set for imaging results using traditional bandpass filters in different frequency ranges (red curve, left), and the curve of $\mathcal{M}$ Function (blue curve, right).
(b) The $F_{1}$ scores on the validation set for imaging using traditional bandpass filtering methods and imaging using the generalized bandpass filter obtained directly from AAM.
Figure 14: Results of bandpass range enumeration experiment and network ablation experiment.
@@
(a) Attention distribution of StreakNet-m and the curve of $\mathcal{M}$ Function.
(b) Attention distribution of StreakNetv2-m and the curve of $\mathcal{M}$ Function.
Figure 15: The curve of Attention distribution of StreakNets and the curve of $\mathcal{M}$ Function.

### v3

@@ Abstract
In this paper, we introduce StreakNet-Arch, a real-time, end-to-end binary-classification framework based on our self-developed Underwater Carrier LiDAR-Radar (UCLR) that embeds Self-Attention and our novel Double Branch Cross Attention (DBC-Attention) to enhance scatter suppression. Under controlled water tank validation conditions, StreakNet-Arch with Self-Attention or DBC-Attention outperforms traditional bandpass filtering and achieves higher $F_{1}$ scores than learning-based MP networks and CNNs at comparable model size and complexity. Real-time benchmarks on an NVIDIA RTX 3060 show a constant Average Imaging Time (54 to 84 ms) regardless of frame count, versus a linear increase (58 to 1,257 ms) for conventional methods. To facilitate further research, we contribute a publicly available streak-tube camera image dataset contains 2,695,168 real-world underwater 3D point cloud data. More importantly, we validate our UCLR system in a South China Sea trial, reaching an error of 46mm for 3D target at 1,000 m depth and 20 m range. Source code and data are available at https://github.com/BestAnHongjun/StreakNet.
@@ I Introduction
Underwater laser imaging signal processing technology is crucial for obtaining underwater images, including 2D gray-scale maps and 3D point clouds images, which has wide applications in ocean exploration, biology [1], surveillance [2], archaeology, unmanned underwater vehicles control [3, 4], etc. In contrast to image processing algorithms for underwater image enhancement [5, 6, 7, 8, 9, 10, 11, 12, 13] or restoration [14, 15, 16], underwater laser imaging signal processing technology can process signals from a more fundamental source, such as streak-tube camera and ICCD camera. This approach enables the achievement of superior spatial resolution and extended detection ranges. However, its effectiveness is significantly hindered by a major challenge: scattering. This phenomenon drastically reduces image clarity and limits imaging range.
To address this, the Underwater Carrier LiDAR-Radar (UCLR) employs a suite of strategies to suppress scattering and achieve long-distance underwater imaging [17, 18, 19, 20]. Specifically, the UCLR’s laser source typically utilizes blue or green light to minimize propagation attenuation in water [21, 22], thereby enhancing detection distance. Additionally, a range-gated detector is employed for the UCLR, which is sensitive only to reflected signals received within a specific time window after the pulse is emitted. More importantly, lasers are modulated into high-frequency pulses to exceed the cut-off frequency of water’s low-pass response [23, 24], effectively suppressing light scattering. Since the frequency is typically high ($\geq$100 MHz), receivers employing high temporal resolution optical detection devices are required, such as nanosecond-resolution ICCD camera [25] or picosecond-resolution streak-tube camera [26, 27, 28]. Underwater laser imaging relies on signal processing algorithms to extract target echoes from the received signal. These algorithms determine the presence and arrival time of the echoes, ultimately reconstructing the image. The processing typically involves two stages: scatter suppression and echo identification. Scatter suppression methods in the UCLR include bandpass filtering [26], adaptive filtering [29, 30, 31, 32, 33, 34], and machine learning-based filtering [35]. The objective is to process a signal containing scatter noise into a suppressed scatter signal. Echo identification methods in UCLR primarily rely on thresholding techniques, including manually set thresholds [26] and adaptive thresholds [36], often coupled with matched filtering approaches [26]. These methods aim to determine the presence of echo signals in the received signal.
@@
Figure 1: Overview of the StreakNet-Arch based UCLR system. (a) Pulses from a sub-nanosecond 500 MHz Q-switch laser (532 nm) are split for $t_{G}$-delayed streak-tube triggering and for scanning water-tank targets at $N$ discrete angles via a motorized turntable. (b) Decoding yields $N$ streak-tube images ($2048\times 2048$), one per angle, where the horizontal axis represents time and the vertical axis represents space. Row $j$ in image $i$ encodes the 30 ns light-intensity echo signal at position $j$ and angle $i$. (c) Given the echo signal and template signal, the StreakNet-Arch-based UCLR system outputs (d) the $(j,i)$ pixel in both the $2048\times N$ grayscale and depth maps.
In this paper, we firstly experimented with employing Self-Attention mechanism networks [38] in the signal processing phase of the self-developed UCLR to improve scatter-resistance. This architecture already SOTA in computer vision [39, 40, 41, 42, 43] and NLP [38, 44, 45], emerges as a powerful universal model. To prevent overfitting and boost generalization across scenes, we provide a template signal alongside each input, guiding the network to learn echo-vs-noise distinctions. We further adapt Self-Attention into a Double Branch Cross Attention (DBC-Attention) mechanism, which our validation experiments show yields higher $F_{1}$ scores than both traditional bandpass filtering and contemporaneous learning-based MP and CNN methods at comparable model size and complexity (Table IV) under controlled water tank environment.
Moreover, by recasting imaging as an end-to-end binary classification, our StreakNet-Arch directly flags echo-containing frames, eliminating the batch-wide pending time of conventional algorithms. On an NVIDIA RTX 3060 GPU, StreakNet-Arch achieves a constant Average Imaging Time (AIT) of 54 to 84 ms across up to 64 frames, whereas traditional methods’ AIT grows linearly from 58 ms to 1,257 ms (Fig. 9, Table III), confirming its real-time advantage.
Given that our input comes from streak-tube camera captures, we name this end-to-end framework StreakNet-Arch. Finally, to validate deep-sea performance, we conducted a South China Sea trial, reaching an error of 46mm for 3D target at 1,000 m depth and 20 m range.
@@
    We enhance the UCLR’s signal processing with Self-Attention networks. Further, we propose DBC-Attention, a groundbreaking variant specifically optimized for underwater imaging tasks. Experimental results under controlled water tank environment conclusively demonstrate DBC-Attention’s superiority over the standard Self-Attention approach.
@@
  5. 5.
    We validated the UCLR system in a deep-sea field experiment in the South China Sea, reaching an error of 46mm for 3D target at 1,000 m depth and 20 m.
@@ III-A StreakNet-Arch
The proposed StreakNet-Arch based self-developed UCLR system (Fig. 1a-d) employs a sub-nanosecond Q-switch laser to generate subcarrier-modulated pulses at a frequency of 500 MHz with 532 nm, 80 mJ. A portion of the generated pulse passes through the beam splitter into a delay device, which can be gated by a delay of $t_{G}$ seconds.  [^1] Range-gated imaging technology, which captures images by controlling the camera shutter delay for a certain period $t_{G}$. This enables the reception of signals within a specific range, mitigating the impact of backscattering on imaging. After the delay, a trigger signal is sent to the control circuit of the streak-tube camera. Simultaneously, another part of the pulse is reflected into the water tank.
@@ III-C Attention Analysis Method
Figure 2: Different perspectives of MP model. (a) From the neuron’s perspective, the input to neuron $j$ is computed by summing the products of all input nodes and their associated weights. (b) From the input node’s perspective, if input node $i$ connects to neuron $j$ with weight $w_{ij}$, then neuron $j$ extracts information proportional to $\|w_{ij}\|$ from node $i$, indicating the attention allocated by neuron $j$ to node $i$.
@@ III-F Loss Function
Figure 3: Schematic diagram of dataset partitioning. Approximately 40% of the data were used for training the network, as shown in (a)(d)(g)(j) for depths of 20 m, 15 m, 13 m, and 10 m, respectively. Around 5% were used for $F_{1}$ evaluation on the validation set, as shown in (b)(e)(h)(k). The full dataset was used for visualizing imaging results, as shown in (c)(f)(i)(l).
@@ III-G Datasets
Figure 4: Experimental setup. (a) The 25 m experimental water tank. (b) The 30 cm diameter experimental target. (c) Prototype of the UCLR system.
@@ IV-A Model training
TABLE II: Model Size and Computational Complexity of
Trained Models.
Model Name Model Size Computational Complexity
StreakNet-s 1.09 M 2.40 GFLOPs
StreakNet-m 2.35 M 5.44 GFLOPs
StreakNet-l 6.24 M 17.19 GFLOPs
StreakNet-x 25.05 M 85.83 GFLOPs
StreakNetv2-s 1.12 M 2.40 GFLOPs
StreakNetv2-m 2.61 M 5.44 GFLOPs
StreakNetv2-l 8.35 M 17.19 GFLOPs
StreakNetv2-x 41.87 M 85.83 GLOPs
MP-s 1.16 M 2.46 GFLOPs
MP-m 2.34 M 4.90 GFLOPs
CNN-s 1.06 M 2.26 GFLOPs
CNN-m 2.08 M 4.36 GFLOPs
Under the StreakNet-Arch, we trained StreakNet with the Self-Attention mechanism as the backbone network and StreakNetv2 with the DWC-Attention mechanism for experiments. For comparison, learning-based methods such as the MP networks [37, 35] and convolutional neural networks (CNN) [56] of different scales were also trained concurrently.
During the training phase, the Stochastic Gradient Descent (SGD) algorithm is used to optimize for 120 epochs, with a base learning rate of $2\times 10^{-6}$ per batch. A cosine annealing learning rate strategy is employed, and the Exponential Moving Average (EMA) method is used. The training was performed on a single NVIDIA RTX 3090 (24G). The details of all trained models are shown in Table II.
@@ IV-B StreakNet-Arch exhibits superior anti-scattering capabilities compared to traditional imaging methods
Evaluation on the validation set demonstrates that both Self-Attention-based StreakNet and DWC-Attention-based StreakNetv2 significantly outperform the bandpass filtering algorithm in terms of $F_{1}$ score. Furthermore, with comparable model sizes and computational complexity (see Table II), models under StreakNet-Arch also achieve superior $F_{1}$ scores compared to learning-based MP and CNN models (Table IV). This demonstrates that the StreakNet-Arch has stronger anti-scattering capabilities compared to traditional algorithms. The imaging results are shown in Fig. 5 and 6.
Figure 5: 2D imaging results at depths of 20 m, 15 m, 13 m, and 10 m for (a)–(d) Bandpass, (e)–(h) MP, (i)–(l) CNN, (m)–(p) StreakNet, and (q)–(t) StreakNetv2.
TABLE III: AIT (ms) for traditional imaging algorithms and StreakNet-Arch algorithms.
@@
Figure 6: 3D imaging results at depths of 20 m, 15 m, 13 m, and 10 m for (a)–(d) Bandpass, (e)–(h) MP, (i)–(l) CNN, (m)–(p) StreakNet, and (q)–(t) StreakNetv2.
TABLE IV: $F_{1}$ scores (%) for traditional imaging methods and StreakNet-Arch imaging methods on the validation set.
Baseline Model StreakNet-s StreakNet-m StreakNet-l StreakNet-x StreakNetv2-s StreakNetv2-m StreakNetv2-l StreakNetv2-x
$F_{1}(\%)$ 86.78 88.23 86.71 85.57 86.92 87.03 86.35 86.33
Bandpass 70.82 +15.96 +17.41 +15.89 +14.75 +16.10 +16.21 +15.53 +15.51
MP-s 86.17 +0.61 +2.06 +0.54 -0.6 +0.75 +0.86 +0.18 +0.16
MP-m 86.56 +0.22 +1.67 +0.15 -0.99 +0.36 +0.47 -0.21 -0.23
CNN-s 84.89 +1.89 +3.34 +1.82 +0.68 +2.03 +2.14 +1.46 +1.44
CNN-m 85.12 +1.66 +3.11 +1.59 +0.45 +1.80 +1.91 +1.23 +1.21
The $F_{1}$ gain of StreakNets is highlighted in red for non-learning-based methods. For learning-based methods, the red highlights indicate the $F_{1}$ gain of StreakNet over MP models and CNNs with comparable parameter counts and computational complexity, while all other entries are shown in gray.
Figure 7: The sequence chart of traditional imaging algorithm. Traditional methods require all streak-tube images before thresholding and imaging can proceed, necessitating a complete wait for all data before any result is available.
$\text{AIT}_{\text{traditional}}=\frac{N+1}{2}t_{m}=\frac{N+1}{2}t_{m}.$ (22)
To measure the real-time imaging capability of the algorithm, we propose an evaluation metric for the AIT, defined as the average time from the input of a streak-tube image to obtaining the corresponding imaging result. If we assume $t_{11}=t_{21}=...=t_{N1}=t_{m}$, the AIT for traditional imaging algorithms is calculated using Eq. 22.
Figure 8: The sequence chart of StreakNet-Arch method. StreakNet-Arch enables immediate imaging from each input without waiting for global information.
$\text{AIT}_{\text{streaknet}}=\frac{1}{N}\sum_{i=1}^{N}t_{i1}=t_{m}.$ (23)
It is evident that the AIT of the traditional imaging algorithm is a linear function with respect to $N$, while the AIT of StreakNet-Arch is a constant. Therefore, theoretically, in practical scenarios where $N$ is large, StreakNet-Arch will have a significant advantage in real-time imaging. To validate this theory, we conducted a comparative experiment. We sequentially input $N$ steak-tube images, where $N$ gradually increases from 1 to 64, and tested the AIT metric for both traditional algorithms and StreakNet-Arch on an NVIDIA RTX 3060 (12G) GPU. Experimental results are depicted in Fig. 9 and Table III, with AIT values measured in milliseconds (ms). The experimental findings indicate that as the number of streak-tube images increases from 2 to 64, the AIT for traditional imaging methods escalates linearly from 58 ms to 1257 ms. In contrast, the AIT for the StreakNet method remains constant within the range of 54 ms to 84 ms.
Figure 9: Curve of AIT (ms) with the changing number $N$ of streak-tube images. It is evident that the AIT of the traditional imaging algorithm is a linear function with respect to $N$, while the AIT of StreakNet-Arch is a constant.
@@ IV-D FD Embedding Layer is an equivalent bandpass filter
Since the carrier frequency of the detection signal is 500 MHz (see Fig. 1e), traditional bandpass imaging algorithms use a handcraft bandpass filter with a range of 450 MHz - 550 MHz during filtering. If we consider the bandpass filter from the perspective of “attention distribution”, we can think of the bandpass filter as a binary attention distribution with values of 1 for frequencies in the range of 450 MHz - 550 MHz and 0 for frequencies outside this range. The FD Embedding Layer’s attention distribution offers a similar concept, functioning as a learnable generalized bandpass filter.
@@
Figure 10: Results of attention distribution with frequency after AAM analysis. (a) Attention distribution with frequency after AAM analysis for StreakNet-m. (b) Attention distribution with frequency after AAM analysis for StreakNetv2-m.
Therefore, we further enumerated the range of bandpass filters in the range of 0 - 200 MHz, with each group spanning 5 MHz, and used traditional bandpass methods for imaging. We then calculated the $F_{1}$ score on the validation set. The experimental results are shown as the red curve in Fig. 11a. A peak appears at 42.5 MHz (i.e., the 40 MHz - 45 MHz bandpass range), indicating that frequency information near 40 MHz is indeed strongly correlated with anti-scattering imaging.
@@
Figure 11: Results of bandpass range enumeration experiment and network ablation experiment. (a) The $F_{1}$ scores on the validation set for imaging results using traditional bandpass filters in different frequency ranges (red curve, left), and the curve of $\mathcal{M}$ Function (blue curve, right). (b) The $F_{1}$ scores on the validation set for imaging using traditional bandpass filtering methods and imaging using the generalized bandpass filter obtained directly from AAM.
@@
Figure 12: The curve of Attention distribution of StreakNets and the curve of $\mathcal{M}$ Function. (a) Attention distribution of StreakNet-m and the curve of $\mathcal{M}$ Function. (b) Attention distribution of StreakNetv2-m and the curve of $\mathcal{M}$ Function.